A store mixes Brazilian coffee worth $11 per kilogram and Venezuelan coffee worth $14 per kilogram. The mixture is to sell for $12 per kilogram. Find how much of each should be used to make a 600-kilogram mixture.
Answer by mananth(16946)

--------price ----------------quantity
Brazilian coffee $ 11---------------- x kgs
Venezuelan coffee $14 ----------------600-x kgs
Mixture $12.00 ---------------- 600

11x+14(600-x)=12*600

11x+8400-14x=7200
11x-14x =7200-8400
-3x=-1200
/-3
x=400 kgs $11.00
200 kgs $14.00